# Maud Biquard ISAE-Supaero / CNES

### v1

@@ 6.3 Results on real satellite images
Figure 7: Handling of compression artefacts for each method. Restoration results of several methods on a real DENVER satellite image of TEST30_REAL, for the IR problem. CI stands for the commercialized image. ©CNES 2024
Figure 8: Results for several methods on a real DENVER satellite image of TEST30_REAL, for IR+SISR problem. CI stands for commercialized image.©CNES 2024
In this section, we present the results of the different methods on TEST30_REAL. As we do not have access to any ground truth, the analysis will be essentially based on visual results. Additionally, we provide BRISQUE scores in Tab. 6, though this metric exhibits high variance and might not be fully adapted to satellite images. CI refers to the Commercialized (Pléiade) Image. Direct inversion methods exhibit the best BRISQUE scores for the IR+SISR problem, which can be explained by their sharp results involving ”typical” high frequencies that are not necessarily accurate. However, in IR, VBLE-xz is competitive with the direct inversion baselines.
@@
IR ATAC. 41.20 51.09 24.85 50.11 46.96 32.17
DENV. 27.48 27.88 30.34 23.89 24.74 21.62
SHAN. 32.43 31.93 27.21 27.90 31.40 27.23
[…]
IR+SISR ATAC. x 56.30 19.13 30.51 54.57 46.08
DENV. x 40.15 5.42 14.39 36.47 33.53
SHAN. x 39.31 5.79 12.64 42.64 35.98
[…]
@@
Visual results are given in Fig. 8 and in Fig. 8. In Fig. 8, the measure contains some compression artefacts which are amplified by in the commercialized image (CI). All deep learning methods perform better than CI for removing these artefacts. However, for RDN and SRResNet, some of the artefacts still slightly remain, while DR methods manage to remove them entirely. Indeed, direct inversion methods yield sharper results but deconvolve details that are not present in the true image.
Figure 8 shows the results of image restoration with super-resolution for all methods. For VBLE-xz, the MMSE estimate, as well as two posterior samples and the predicted 90% perpixel quantiles are provided. The zoomed commercialized image is very blurry, direct inversion methods yield sharp results, but with some artefacts, while DR methods provide smoother results. SatDPIR provides excellent results, however, it sometimes yields lines that do not exist. VBLE-xz provides very consistent results, its MMSE estimate is rather smooth, but the samples are sharper and contain valuable information. For instance, in the zoomed area, two forms are proposed for the linear and both seem plausible. Besides, the 90% quantile map provides a meaningful way to localize unsure areas.
@@ 6.4 Robustness evaluation
In this section, we evaluate the robustness of the different methods to resolution and model changes using TEST30_SIMU.
6.4.1 Resolution robustness
Direct inversion models are problem specific, therefore need to be retrained or fine-tuned for each resolution and each inverse problem. For DR methods VBLE-xz and SatDPIR, the models can be used to restore several inverse problems, yet, it should a priori be retrained for each resolution. In the resolution experiments, we evaluate the robustness on IR and IR+SISR problems. For IR, we use the DR models trained on 25cm target images and evaluate them on 50cm images. For direct inversion methods, RDN and SRResNet are trained to deblur 25cm images and are evaluated on 50cm images. For IR+SISR, we use the DR models trained at 50cm and evaluate them on 25cm images.
@@ 6.4.1 Resolution robustness
Quantitative results are provided in Tab. 7. Surprisingly, for all methods, the performance is not significantly affected by the resolution change. This could be explained by the fact that some invariance to resolution exists in several satellite landscapes, for instance in textured areas. It is however not the case for some small or high frequency structures, such as cars or pedestrian crossings. In Fig. 9, which provides the restoration of a pedestrian crossing for each method at the two resolutions, the crossing is restored worse with the model trained at the wrong resolution. This is particularly visible for SatDPIR as it restores particularly well the crossing at the right resolution.
@@
Figure 9: Restored image from TEST30_SIMU by several methods for the IR problem, using a network trained at the right resolution (first row), and a network trained at the wrong resolution (second row). ©CNES 2024
6.4.2 Robustness to model changes
@@
Then, we evaluate the robustness of the different methods to changes in the forward model. We choose, for simplicity, to induce slight changes in the PSF, parameterized by the Modulated Fourier Function (MTF), that is the Fourier transform of the PSF, evaluated at $f_{e}/2$. The Pléiades MTF at $f_{e}/2$ is $0.15$. We compute image restoration results for MTF values of $0.13$ and $0.12$ at $f_{e}/2$. These values are representative of the evolution of the instrument PSF in orbit. As direct inversion methods are trained for a specific PSF with MTF $=0.15$, they should not be very robust to these changes. For DR methods, if the change of PSF is taken into account in the forward model during the restoration, the performance should not be affected. A visual illustration is given in Fig. 11, while quantitative results are provided in Fig. 10.
@@
Figure 10: PSNR and LPIPS of several methods with changes in the forward model. Dashed lines denote the results without modelling the change. For DR methods SatDPIR and VBLE, continuous lines denote the results when the change is modelled in the forward model during restoration. MTF=0.15 at $f_{e}/2$ is the original value of the MTF in the experiments.
@@
Figure 11: Restored image from TEST30_SIMU by several methods for the IR problem with changes in the degradation model (MTF=0.12 at $f_{e}/2$). This forward model change is not modelled, except for DR methods when True MTF is mentioned. ©CNES 2024
@@
Dashed lines in the figure represent the performance of the methods when the wrong PSF, with MTF$=0.15$, is used for image restoration. Continuous lines for DR methods represent the performance when the PSF change is integrated into the forward model. When the wrong PSF is used for DR methods, the performance decay is at the same rate as RDN and SRResNet methods. Yet, when the PSF change is integrated in the forward model, the performance is not affected, showing that DR methods can be used with different PSFs without any retraining, as long as the forward model is correctly specified. This can be crucial in practice as the PSF of a satellite evolves with time, mainly due to focus changes. Conversely, direct inversion networks should be retrained each time the PSF is significantly modified.
@@
Therefore, SatDPIR and VBLE-xz are robust to resolution changes and to changes in the forward model. Additionally, they also work very well on real data, as shown in the previous section. These data contain new types of landscapes that have not been seen during training, i.e. clouds and Atacama desert, and additional model errors induced by compression and satellite off-pointing. This shows that the methods are robust to a wide range of problems and can be used in practice in real satellite processing chains.
@@ 6.5 Ablation study
DPIR SatDPIR
IR PSNR $\uparrow$ 48.70 48.66
LPIPS $\downarrow$ 0.0139 0.0138
Time 4.4s 3.8s
IR+SISR PSNR $\uparrow$ 37.19 37.18
LPIPS $\downarrow$ 0.1656 0.1658
Time 18min09s 1min08s
Table 8: Comparison of DPIR and our proposed method SatDPIR. Results on TEST30_SIMU.
Table 8 highlights the interest of our proposed method, SatDPIR, compared to DPIR. SatDPIR has almost exactly the same metrics as DPIR, thus the solutions found by SatDPIR and DPIR are very close. Yet, SatDPIR is much faster especially when super-resolution is performed jointly to restoration. Indeed, for this problem, the computation of the data-fidelity proximal operator is very slow and the proposed accelerated framework enables to get results 20 times faster than DPIR.
Figure 12: PSNR, LPIPS and 90% ICP of VBLE-xz (not calibrated) on TEST14_SIMU for the IR problem, for varying regularization parameter $\lambda$.
Figure 12 shows diverse metrics for VBLE-xz for different values of $\lambda$ regularization parameter. $\lambda=1$ corresponds to the exact Bayesian framework. While the best PSNR and LPIPS are reached around $\lambda=0.4$, the quality of the predicted (uncalibrated) ICPs strongly decreases for $\lambda<0.6$. Hence, we have chosen $\lambda=0.6$ for all the experiments, as it seems a good trade-off between the different metrics.
@@ 7 Conclusion
In this paper, we have introduced the VBLE-xz method for satellite image restoration, which estimates the inverse problem posterior using variational inference by capturing the uncertainty jointly in the latent and image space of a variational compressive autoencoder. We also proposed SatDPIR, which accelerates the original plug-and-play framework which, in the case of Poisson-Gaussian noise, requires an inner iteration process.
The conducted experiments demonstrate the interest of VBLE-xz as a scalable posterior sampling method for satellite image restoration, as it outperforms the state-of-the-art PnP-ULA method both in terms of performance and posterior samples realism while being by several orders of magnitude faster. Furthermore, the results of VBLE-xz and SatDPIR visually exhibit very few artefacts compared to state-of-the-art direct inversion methods, which is crucial for current satellite super-resolution chains. Quantitatively, VBLE-xz remains competitive with these baselines for point estimation results, and SatDPIR outperforms them in terms of PSNR and SSIM, excelling in deconvolving small details. This makes SatDPIR a very interesting alternative to direct inversion methods if only point estimation is required.
Future works will be dedicated to improving and extending the method, for instance by considering variable bitrate and fixed-quality CAEs [32], that would enable a local tuning of the regularization and should improve the results when the image contains both high-frequency and smooth areas. Another interesting perspective will be to demonstrate the interest in the generated uncertainties for real remote sensing use cases such as change detection or land-cover mapping.

### v2

@@ 5.3 Results on real satellite images
Figure 7: Handling of compression artifacts for each method. Restoration results of several methods on a real DENVER satellite image of TEST30_REAL, for the IR problem. CI stands for the commercialized image. ©CNES 2025
Figure 8: Results for several methods on a real DENVER satellite image of TEST30_REAL, for IR+SISR problem. CI stands for commercialized image.©CNES 2025
In this section, we present the results of the different methods on TEST30_REAL. As we do not have access to any ground truth, the analysis will be essentially based on visual results. Additionally, we provide BRISQUE scores in Tab. 5, though this metric exhibits high variance and might not be fully adapted to satellite images. CI refers to the currently Commercialized (Pléiade) Image. Direct inversion methods exhibit the best BRISQUE scores for the IR+SISR problem, which can be explained by their sharp results involving ”typical” high frequencies that are not necessarily accurate. However, in IR, VBLE-xz is competitive with the direct inversion baselines.
@@
IR ATAC. 41.20 51.09 24.85 50.11 46.96 32.27
DENV. 27.48 27.88 30.34 23.89 24.74 20.81
SHAN. 32.43 31.93 27.21 27.90 31.40 26.19
TOTAL 33.70 36.97 24.13 33.97 34.37 26.42
IR+SISR ATAC. x 56.30 19.13 30.51 54.57 36.39
DENV. x 40.15 5.42 14.39 36.47 40.16
SHAN. x 39.31 5.79 12.64 42.64 40.15
TOTAL x 45.25 10.11 19.18 44.56 38.90
@@
Visual results are given in Fig. 8 and in Fig. 8. In Fig. 8, the measure contains some compression artifacts which are amplified by the restoration performed in the commercialized image (CI). All deep learning methods perform better than CI for removing these artifacts. However, for direct inversion methods RDN and SRResNet, some of the artifacts still slightly remain, while DR methods manage to remove them entirely. Indeed, direct inversion methods yield sharper results but deconvolve details that do not lie in the true image.
Figure 8 shows the results of image restoration with super-resolution for all methods. For VBLE-xz, the MMSE estimate, as well as two posterior samples and the predicted 90% perpixel quantiles are provided. The zoomed commercialized image is very blurry, direct inversion methods yield sharp results, but with some artifacts, while DR methods yield smoother results. PG-DPIR provides excellent results, however, it sometimes yields lines that do not exist. VBLE-xz exhibits very consistent results, its MMSE estimate is smoother than images obtained by direct inversion, but the samples are sharper and contain valuable information, for instance concerning the shapes of the linears. Besides, the 90% quantile map provides a meaningful way to localize unsure areas.
@@ 5.4 Robustness evaluation
In this section, we evaluate the robustness of the different methods to resolution and modeling errors using TEST30.
5.4.1 Robustness to resolution change
Direct inversion models are problem-specific, therefore they need to be retrained or fine-tuned for each resolution and each inverse problem. For DR methods VBLE-xz and PG-DPIR, the models can be used to restore several inverse problems, yet, it should a priori be retrained for each resolution. In the following experiments, we evaluate the robustness to resolution change for IR and IR+SISR problems. For IR, we use the DR models trained on 25cm target images and evaluate them on 50cm images. For direct inversion methods, RDN and SRResNet are trained to deblur 25cm images and are evaluated on 50cm images. For IR+SISR, we use the DR models trained at 50cm and evaluate them on 25cm images.
@@ 5.4.1 Robustness to resolution change
Quantitative results are provided in Tab. 6. Surprisingly, for all methods, the performance is not significantly affected by the resolution change. This could be explained by the fact that some invariance to resolution exists in several satellite landscapes, for instance in textured areas. However, it is not the case for some small or high-frequency structures, such as cars or pedestrian crossings. In Fig. 9, which provides the restoration of a pedestrian crossing for each method at the two resolutions, the crossing is restored worse with the model trained at the wrong resolution. This is particularly visible for PG-DPIR as it restores particularly well the crossing at the right resolution.
@@
Figure 9: Restored image from TEST30 by several methods for the IR problem, using a network trained at the right resolution (first row), and a network trained at the wrong resolution (second row). ©CNES 2025
5.4.2 Robustness to modeling errors
@@
Here, we evaluate the robustness of the different methods to modeling errors. Note that testing on real images a model trained only on simulated images is a disadvantageous scenario for performance assessment. However, it should be emphasized that it affects differently DR and direct inversion methods. Indeed, simulating realistic target images is quite feasible when very high-resolution airborne data are available. On the opposite, generating realistic degraded images is much more challenging. This difficulty arises because not all the physical phenomena can be accurately modeled. For instance, the blur is acquisition-dependent and varies with the satellite orientation. As a result, a domain gap exists when training direct inversion methods with pairs of simulated target and degraded images. DR methods, by contrast, do not seem to suffer from this domain gap since they are trained solely on target images. However, note that DR methods are impacted by the modeling errors regarding the satellite forward model during restoration step. As quantifying the performance gap occurring when restoring real data is hardly feasible as no ground truth data are available, we choose, for simplicity, to induce slight changes in the PSF, parameterized by the Modulated Transfer Function (MTF), that is the Fourier transform of the PSF, evaluated at $f_{e}/2$. The Pléiades MTF at $f_{e}/2$ is $0.15$. We compute image restoration results for MTF values of $0.13$ and $0.12$ at $f_{e}/2$. These values are representative of the evolution of the instrument PSF in orbit. As direct inversion methods are trained for a specific PSF with MTF =0.15=0.15, they should not be very robust to these changes. For DR methods, if the change of PSF is taken into account in the forward model during the restoration, the performance should not be affected. A visual illustration is given in Fig. 11, while quantitative results are provided in Fig. 10.
@@
Figure 10: PSNR and LPIPS of several methods with changes in the forward model. Dashed lines denote the results without modeling the change. For DR methods PG-DPIR and VBLE-xz, continuous lines denote the results when the change is modeled in the forward model during restoration. MTF=0.15 at $f_{e}/2$ is the original value of the MTF in the experiments.
@@
Figure 11: Restored image from TEST30 by several methods for the IR problem with changes in the degradation model (MTF=0.12 at $f_{e}/2$). This forward model change is not modeled, except for DR methods when True MTF is mentioned. ©CNES 2025
@@
Dashed lines in the figure represent the performance of the methods when the wrong PSF, with MTF=0.15=0.15, is used for image restoration. Continuous lines for DR methods represent the performance when the PSF change is integrated into the forward model. When the wrong PSF is used for DR methods, the performance decay is at the same rate as for RDN and SRResNet methods. Yet, when the PSF change is integrated into the forward model, the performance is not affected, showing that DR methods can be used with different PSFs without any retraining, as long as the forward model is correctly specified. This can be crucial in practice as the PSF of a satellite evolves with time, mainly due to focus changes. Conversely, direct inversion networks should be retrained each time the PSF is significantly modified.
@@
Therefore, PG-DPIR and VBLE-xz are robust to changes in the resolution and in the forward model. Additionally, they also work very well on real data, as shown in the previous section. These data contain new types of landscapes that have not been seen during training, i.e. clouds and Atacama desert, and additional model errors induced by compression and satellite off-pointing. This shows that the methods are robust to a wide range of problems and can be used in practice in real satellite processing chains.
@@ 5.5 Ablation study
Figure 12: Rate-distortion curve computed on TEST30. Blue curve: each point corresponds to a CAE trained at a specific bitrate. Orange curve: One CAE trained at a high bitrate, with different multirate scaling factors.
In this section, we demonstrate the usefulness of the additional scaling factor $s$, which enables the modulation of the CAE bitrate, and thus of the regularization strength when solving different inverse problems. Figure 12 shows two rate-distortion curves obtained on TEST30, illustrating the trade-off between compression efficiency and reconstruction error for a given compressor. The blue curve corresponds to the rate-distortion trade-offs of several single-rate CAEs, each trained with a different $\alpha$, as defined in eq. 7. The orange curve represents the trade-offs obtained using the CAE trained for the highest bitrate, while varying the scaling factor $s$ to modulate the rate, following the approach described in [Arezki2024] and applied in VBLE-xz. For $s\in[0.2,1]$, the two curves are nearly identical, validating the effectiveness of this multirate approach for satellite data.
PSNR SSIM LPIPS ICP 90%
CAE $\alpha=0.0932,s=1$ 37.35 0.9490 0.1698 0.87
CAE $\alpha=0.36,s=0.4$ 37.31 0.9484 0.1691 0.88
Table 7: Results of IR+SISR on TEST14. $\alpha$ is the rate-distortion parameter of eq. 7, a high $\alpha$ corresponds to a high bitrate. $s$ is the multirate scaling factor.
Additionally, some image restoration results on the IR+SISR problem are provided in Tab. 7, comparing the use of a CAE trained at an adapted low bitrate with $s=1$ and a CAE trained at a high bitrate with $s=0.4$. Both networks yield very similar performance. This demonstrates the relevance of incorporating the scaling factor $s$ in VBLE-xz. Indeed, it enables control over the regularization through a single additional hyperparameter, rather than requiring the training of another CAE with an appropriate rate-distortion trade-off.
@@ 6 Conclusion
In this paper, we proposed the VBLE-xz method for satellite image restoration. VBLE-xz estimates the inverse problem posterior using variational inference by capturing the uncertainty jointly in the latent and the image space of a variational CAE. We also proposed to use a scaling factor to modulate the CAE bitrate, a remarkably simple way to adapt the regularization to a given inverse problem difficulty.
While deep regularization (DR) methods possess various advantages compared to direct inversion methods - in particular restoring images from different sources with a single network, and avoiding the use of simulated degraded images for training -, our experiments demonstrate that the DR methods VBLE-xz and PG-DPIR also present advantages in terms of restoration performance. In particular, they provide consistent performance at a reasonable computational cost and introduce fewer artifacts than direct inversion. When uncertainty quantification (UQ) is required, VBLE-xz offers the best trade-off between image quality, UQ accuracy, and computation time. In contrast, when UQ is not needed, PG-DPIR achieves slightly better performance, yet with more pronounced artifacts.
Future works will be dedicated to improving and extending the method, for instance, by introducing correlations in the image space uncertainty model. Another interesting perspective will be to demonstrate the interest of the generated uncertainties for real remote sensing use cases such as change detection or land-cover mapping.